# Burnside type results for fusion categories

### v10

@@ 9. Applications to fusion categories
Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring. As in §1, we denote the set of isomorphism classes of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$. Let also $d_{i}:=\dim(X_{i})$ be the quantum dimension of $X_{i}$ for all $i\in\mathcal{I}$. For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
Let ${\mathcal{Z}(\mathcal{C})}$ be the Drinfeld center of the fusion category $\mathcal{C}$. The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$. It is known that $A:=R(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{Z}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $A$. The vector space $\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)$ is called the set of central elements. The space ${\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$ is called the space of class functions of $\mathcal{C}$. Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character, [Shi17]. The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
@@
where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$. Since ${\mathcal{Z}(\mathcal{C})}$ is also fusion category we can write $R(\mathbf{1})=\bigoplus_{j=1}^{m}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$. Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$. The above isomorphism also gives a canonical bijection between the set of central primitive idempotents $\{F_{j}\}_{j=1}^{m}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$. We denote as above $\mathcal{J}:=\{1,\dots m\}$.
For any fusion category $\mathcal{C}$, by abuse of notation, the symmetric RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$. We denote by ${\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
@@
in the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$. In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore ${\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
@@
Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let ${\widehat{{K(\mathcal{C})}}}$ be its dual. Then
(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix determined by $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
@@
Note that the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21]. Theorem 3.4 from the same paper implies that for any pseudo-unitary fusion category there is a canonical isomorphism of hypergroups
(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})}.$
For a pivotal fusion category, Shimizu has defined $C_{j}:={\mathcal{F}_{\lambda}}^{-1}(F_{j})\in\mathrm{{CE}(\mathcal{C})}$ to be the conjugacy class sums corresponding to the conjugacy class $\mathcal{C}^{j}$. Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5]. Recall also the Fourier transform of $\mathcal{C}$ associated to $\lambda$ is the linear map
@@
Since any weakly-integral fusion category is pseudo-unitary, by applying the canonical morphism $\alpha$ one obtains the following:
@@
For any weakly-internal fusion category $\mathcal{C}$ the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
(9.6) $\prod_{j=1}^{m}C_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
@@
Let $\mathcal{C}$ be a spherical fusion category with the Grothendieck ring ${K(\mathcal{C})}$ commutative. Let also $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ a character of ${K(\mathcal{C})}$. Suppose that ${\widehat{{K(\mathcal{C})}}}$ is a RN-hypergroup.
Then $\mu_{j}$ is a grouplike element of ${\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
@@
By Lemma 2.17, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}d_{j}^{2}=1$. Since ${\widehat{{K(\mathcal{C})}}}$ is normalized it follows that $d_{j}=1$ for all $j\in\mathcal{J}$. Thus $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$. On the other hand, it follows from Equation (9.2) that ${\widehat{h}}_{j}=1$ if and only if $\mathcal{C}^{j}$ is a one dimensional conjugacy class of $\mathcal{C}$. ∎
Suppose now that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. There is a subset $\mathcal{J}_{\mathcal{D}}\subseteq\mathcal{J}$ such that
(9.7) $\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{J}_{\mathcal{D}}}F_{j}$
since $\lambda_{\mathcal{D}}$ is an idempotent element inside ${\mathrm{CF}(\mathcal{C})}$. Note that this set is denoted by $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Also, this is the corresponding notion of the support from Definition 6.3.
Note that [Bur20, Lemma 4.6 ] implies
(9.8) $J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{J}\;|\;\dim(\mathcal{C}^{j})=1\},$
@@
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
(9.9) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in J_{{\mathcal{C}_{\mathrm{ad}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
@@
Suppose that the dual ${\widehat{{K(\mathcal{C})}}}$ is Burnside. Squaring Equation (9.9) it follows that
$\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\sum_{j\in J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
The converse also follows immediately from Proposition 9.5. ∎
Let us provide a dual statement in the group case:
@@
For any finite nilpotent group $G$ the dual $\widehat{K(G)}$ is Burnside and therefore
@@
Let $\mathcal{C}=\mathrm{Rep}(G)$. It is easy to verify that $\widehat{K(G)}$ is a real-non negative hypergroup. For, example see [Bur21, Equation (3.12)]. Then Theorem B of [INW99] says that if $G$ is a nilpotent group then ${K(G)}$ is dual-Burnside. ∎
Theorem 1.12 implies the following:
@@
Suppose that $\mathcal{C}$ is a nilpotent fusion category. Then ${K(\mathcal{C})}$ and ${\widehat{{K(\mathcal{C})}}}$ are both Burnside.
@@ 10. Premodular categories
Recall that a premodular category is a braided and spherical fusion category. For the rest of this section suppose that $\mathcal{C}$ is a pseudo-unitary premodular category. By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{i^{\prime}=1}^{m}\frac{s_{ii^{\prime}}}{d_{i^{\prime}}}E_{i^{\prime}}.$
where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $E_{i^{\prime}}$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
As in [Bur22, §4] there is a braided function $M:\mathcal{I}\rightarrow\mathcal{J}$ such that if
$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i}$
[…]
Recall from [BB15] the notion of cosets of a fusion category with respect to a fusion subcategory. By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$. For any $j\in\mathcal{J}_{2}$ we denote by $\mathcal{R}_{j}:=\{[X_{i}]\;|\;M(i)=j\}$. The aforementioned theorem implies also that $\mathcal{R}_{j}$ with $j\in\mathcal{J}_{2}$ are exactly the right cosets of $\mathcal{C}$ with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$. Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters. By [Bur22, Equation (4.20)],
(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{J}_{2}.$
@@
Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{Z}_{2}(\mathcal{C})\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and it acts freely on the set $\mathrm{Irr}(\mathcal{C})$ of isomorphism classes of simple objects of $\mathcal{C}$. Then for any $j\in\mathcal{J}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
@@
Note that in ${K(\mathcal{C})}$ one has $h_{i}=1$ for all $i$, and therefore ${K(\mathcal{C})}$ is $h$-integral. If $\mathcal{C}$ is weakly-integral, the fact that the dual ${\widehat{{K(\mathcal{C})}}}$ is rational RN follows from [Bur21, Equation (4.5)]. Thus one may try to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$. In order to do be able to do this, one needs to verify that the additional condition (4.18) is satisfied for all $1\leq i\leq m$, i.e. $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$. In [Bur22, Lemma 4.2] it was shown that
@@
for all $1\leq i,i^{\prime}\leq m$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j$. It follows that
@@
Since $\mathcal{C}$ is spherical one has $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
(10.6) $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}_{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
We will show that $\dim(\mathcal{C}^{j})=d_{i_{j}}^{2}$ and then the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$. Indeed, applying Equation (10.2), then $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$ since the action of $\mathcal{C}^{\prime}$ is free on $\mathrm{Irr}(\mathcal{C})$. ∎

### v11

@@ 9. Applications to fusion categories
Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring. As in §1, we denote the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$. Let also $d_{i}:=\dim(X_{i})$ be the categorical dimension of $X_{i}$ for all $i\in\mathcal{I}$. For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
Let ${\mathcal{Z}(\mathcal{C})}$ be the Drinfeld center of the fusion category $\mathcal{C}$. The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$. It is well-known that $A:=FR(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{C}$ (meaning that $R(\mathbf{1})$ is a commutative algebra in ${\mathcal{Z}(\mathcal{C})}$), where $\mathbf{1}$ is the unit object of $\mathcal{C}$. The vector spaces
$\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)\text{ and }{\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$
are respectively called the space of central elements and the space of class functions of $\mathcal{C}$. Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})_{\mathbb{C}}$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character [Shi17, Definition 3.9]. The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
@@
where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$. Since ${\mathcal{Z}(\mathcal{C})}$ is also a fusion category we can write $R(\mathbf{1})=\bigoplus_{j\in\mathcal{I}}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$. Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$. The above isomorphism also gives a canonical bijection between the set of primitive central idempotents $\{F_{j}\}_{j\in\mathcal{I}}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$.
For any fusion category $\mathcal{C}$, by abuse of notation, the RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$. We denote by $^{\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
@@
in the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$. In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
@@
Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let $^{\widehat{{K(\mathcal{C})}}}$ be its dual. Then
(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i}$
where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
@@
Note that the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21]. Theorem 3.4 from the same paper implies that for any pivotal fusion category there is a canonical isomorphism of $\mathbb{C}$-algebra
(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})},$
where $C_{j}\in\mathrm{{CE}(\mathcal{C})}$ is the conjugacy class sum corresponding to the conjugacy class $\mathcal{C}^{j}$, defined for a pivotal fusion category by Shimizu as ${\mathcal{F}}^{-1}(F_{j})$. Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5]. Recall also the Fourier transform of $\mathcal{C}$ associated to $\lambda$ is the linear map
@@
By [Bur20, Lemma 4.1 and (4.7)], for any spherical fusion category $\mathcal{C}$ over $\mathbb{C}$ with a commutative Grothendieck ring, we have
$\langle F_{i},C_{j}\rangle=\delta_{i,j}\dim(\mathcal{C}^{j}).$
This relation implies that the set $(F_{i})$ forms a dual basis to $(C_{j}/\dim(\mathcal{C}^{j}))$ with respect to the bilinear form $\langle$,$\rangle$ , ⟩. Consequently, this provides an easier way for defining the conjugacy class sums $(C_{j})$ in this case (see [Bur20, (4.11)] for a formula).
@@
For any weakly-integral fusion category $\mathcal{C}$ with a commutative Grothendieck ring the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
(9.6) $\prod_{j=1}^{m}\frac{C_{j}}{\dim(\mathcal{C}^{j})}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
@@
Proof.
By Corollary 4.21, we can apply Corollary 9.2. The result follows by applying the canonical isomorphism $\alpha$. ∎
@@
Let $\mathcal{C}$ be a spherical fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$ and RN dual. Let $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ be a linear character. Then $\mu_{j}$ is a grouplike element of $^{\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
@@
By Lemma 2.24 and the fact that $^{\widehat{{K(\mathcal{C})}}}$ is normalized, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$. This is also equivalent to $\dim(\mathcal{C}^{j})=1$, by Equation (9.2). ∎
Now, suppose that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. Following the notion of support defined in Definition 6.3, there exists a subset $\mathcal{I}_{\mathcal{D}}\subseteq\mathcal{I}$ such that
$\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{I}_{\mathcal{D}}}F_{j},$
Please note that this set is called $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Additionally, [Bur20, Lemma 4.6] implies
(9.7) ${\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{I}\;|\;\dim(\mathcal{C}^{j})=1\},$
@@
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then $^{\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
(9.8) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
@@
It follows from Corollary 4.9 and Theorem (6.5), together with
$\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)},$
where the last equality comes from Equation (5.2). ∎
Here is a reformulation of Theorem 1.12 for the case of a nilpotent finite group:
@@
For any finite nilpotent group $G$, the dual $\widehat{\text{ch}(G)}$ is Burnside and therefore
@@
Recall that $K(\text{Rep}(G))$ is the character ring $\text{ch}(G)$. The result that $\widehat{\text{ch}(G)}$ is a RN hypergroup is established in [Bur21, Equation (3.12)]. Furthermore, [INW99, Theorem B] states that if $G$ is a nilpotent group, then $\text{ch}(G)$ is dual-Burnside. ∎
Theorem 1.15 leads to the following conclusion:
@@
If $\mathcal{C}$ is a nilpotent fusion category with a commutative Grothendieck ring and RN dual, then both ${K(\mathcal{C})}$ and $^{\widehat{{K(\mathcal{C})}}}$ are Burnside.
Remark 9.8.
A nilpotent fusion category is weakly-integral [GN08], which means it is also pseudo-unitary and spherical [EGNO15]. It remains an open question whether every pseudo-unitary fusion category can be given a unitary structure [EGNO15, Remark 9.4.7]. However, a commutative fusion ring with a unitary categorification does have a RN dual [LPW21]. Therefore, we believe that the assumption of a RN dual in Corollary 9.7 could be omitted. Additionally, in the braided case, its Grothendieck ring possesses a RN dual according to [Bur21, Theorem 1.2].
@@ 10. Premodular categories
Recall that a premodular category is defined as a braided spherical fusion category. In the rest of this section, we will assume that $\mathcal{C}$ is a premodular category. By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{j\in\mathcal{I}}\frac{s_{ij}}{d_{j}}E_{j}.$
where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $(E_{j})$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
As described in [Bur22, §4], there is a map $M:\mathcal{I}\to\mathcal{I}$ (referred to as the braided partition function) such that if
$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i},$
then $M(i)=j$. Let $\mathcal{I}_{2}\subseteq\mathcal{I}$ denote the set of all indices $j$ for which $\mathrm{f}_{Q}(F_{j})\neq 0$, meaning that $\mathcal{B}_{j}$ is not an empty set. Given that $\mathrm{f}_{Q}(1)=1$, we can thus partition the set of all isomorphism classes of simple objects $\text{Irr}(\mathcal{C})=\bigsqcup_{j\in\mathcal{I}_{2}}\mathcal{R}_{j}$, where $\mathcal{R}_{j}=\{[X_{i}]\;|\;i\in\mathcal{B}_{j}\}$. In other words, $\mathcal{R}_{j}=\{[X_{i}]\;|\;M(i)=j\}$. Thus, we obtain a unique function $M:\mathcal{I}\to\mathcal{I}_{2}$ with the property that $E_{i}\mathrm{f}_{Q}(F_{M(i)})\neq 0$ for all $i\in\mathcal{I}$.
The paper [BB15] introduces the notion of cosets of a fusion category $\mathcal{C}$ with respect to a fusion subcategory $\mathcal{D}$: two simple objects $X,Y$ in $\mathcal{C}$ are in the same (right) coset if and only if there is a simple object $S$ in $\mathcal{D}$ such that $X$ is a constituent of $Y\otimes S$. By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$. In other words, $(\mathcal{R}_{j})_{j\in\mathcal{I}_{2}}$ correspond exactly to these cosets. Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters. By [Bur22, Equation (4.20)],
(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{I}_{2}.$
@@
Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and acting freely on the set $\mathrm{Irr}(\mathcal{C})$. Then for any $j\in\mathcal{I}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
@@
First of all, ${K(\mathcal{C})}$ is $h$-integral because as a fusion ring, $h_{i}=1$ for all $i$. If $\mathcal{C}$ is weakly integral, then the dual $^{\widehat{{K(\mathcal{C})}}}$ is a rational RN hypergroup, as shown by [Bur21, Equation (4.3)]. Therefore, we can attempt to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$. To do this, we need to ensure that the additional condition (4.18) holds for all $i\in\mathcal{I}$, specifically that $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$, where ${\alpha_{{}_{ij}}}:=\mu_{j}(x_{i})$. It was demonstrated in [Bur22, Lemma 4.2] that
@@
for all $i,i^{\prime}\in\mathcal{I}$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j\in\mathcal{I}_{2}$, and take $i^{\prime}=i_{j}$. It follows that
@@
But $\mathcal{C}$ is spherical, so $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
(10.6) $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}^{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
But $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$, acting freely on $\mathrm{Irr}(\mathcal{C})$, thus $|\mathcal{R}_{j}|=\dim(\mathcal{C}^{\prime})$, and so by Equation (10.2), $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$, and the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$. ∎
Roughly speaking, Proposition 10.1 can be intuitively understood as follows: a weakly integral premodular fusion category $\mathcal{C}$ is dual-Burnside with respect to $\mathcal{C}^{\prime}$.